OpenStudy (anonymous):

Kary can finish her math final in 3hours. It takes Barry 4 hours to finish the same test. If they work together, how long will it take them to finish the test?

OpenStudy (wolf1728):

If this is the typical "work" problem, the formula is time1 * time2 / (time1 + time2) = total time 3 * 4 / (3 + 4) 12 / 7 = 1 + 5/7 hours = 1 hour 42 minutes 51.43 seconds

OpenStudy (anonymous):

thank you

OpenStudy (wolf1728):

u r welcome

OpenStudy (whpalmer4):

In general, find the rate of each worker: Kary does 1 test/3 hours = 1/3 test/hour Barry does 1 test/4 hours = 1/4 test/hour together they do 1/3+1/4 test/hour = 4/12 + 3/12 = 7/12 test/hour 1 test * 12/7 hour/test = 12/7 hour

OpenStudy (whpalmer4):

rather than dividing by the fraction, I multiplied by its reciprocal, which is equivalent.

OpenStudy (whpalmer4):

Say they got some guy named Bill who is really fast, and does the test in 2 hours: Bill = 1 test/2 hours total rate = 1/2 + 1/3 + 1/4 = 13/12 test/hour time for the bunch to do the test = 1 test * 12/13 hour/test = 55.38 minutes